# Incongruent Melting and Phase Diagram of SiC from Machine Learning Molecular Dynamics

### v1

@@ 3.3 Lower temperature transition from crystalline to decomposed Si + C
Figure 5: Analysis of two-phase coexistence simulations. Panels (left to right) show: fraction of zinc-blende (rock-salt) structures, largest C cluster size (number of atoms), temperature convergence at the NPH stage, over simulation time. Top: 30 GPa. Bottom: 90 GPa.
@@
At the fixed pressure of 30 GPa, during the heating stage (0.5-1 ns), the heated half of the supercell becomes amorphous, and the overall fraction of zinc-blende local structures in the supercell decreases to 0.5. The C clusters begin to form during the cooling stage (1-1.5 ns) and continue growing in size at 1.5-2 ns as demonstrated in the second column of Fig.5. In the NPH simulation stage (1.5-2 ns), both B3 crystal and decomposed Si + C phases coexist, with relatively different equilibrated fractions dependent on the initial temperature. Meanwhile, multiple simulations with different initial temperatures converge to approximately 3500 K during the NPH stage, which is identified as the transition temperature (shown in the third column of Fig.5).
At 90 GPa, where the stable phase at low temperature is rock salt (B1), a similar procedure starting from B1 phase yields a transition temperature around 4500 K. Notably, the carbon cluster sizes formed at 90 GPa are much smaller than those at 30 GPa, leading us to conjecture that at sufficiently high pressures, the decomposed Si+C phase between the crystalline and liquid phases may not be present. Additional coexistence simulations were performed at 10 GPa and 60 GPa, with results presented in the Supplementary Materials.
@@ 3.4 Phase diagram
Here we combine the results obtained from the MD simulations of the SiC phase transitions into a complete pressure-temperature phase diagram. Fig.6 presents the $P$-$T$ diagram, illustrating the B3 $\leftrightarrow$ B1, B3 $\leftrightarrow$ gas, B3 $\leftrightarrow$ Si + C, B1 $\leftrightarrow$ Si + C, and Si + C $\leftrightarrow$ SiC (liquid) transitions. For completeness, we briefly describe the simulations of B3 $\leftrightarrow$ B1 transition and B3 $\leftrightarrow$ gas below, although these are not the primary focus of this work.
@@
The B3 $\leftrightarrow$ B1 transition pressures at various temperatures are determined using the same methods as in our previous work [25]. Starting with the B3 crystal, we increase pressure at a fixed temperature to generate a configuration containing both phases. Then, we perform constant-pressure MD simulations on these coexistence structures at different pressures to identify the transition point, characterized by the stable coexistence of two phases remains. The pressure scan results are shown in the Supplementary Materials. The B3 $\leftrightarrow$ B1 transition pressures identified from our simulations (black solid line in Fig. 6) are close to experimental measurements with a deviation of less than 20 GPa.
At zero or low pressure, the SiC B3 crystal sublimates into gas phase instead of melting, which only occurs under high pressure. Our MD simulation at $T$ = 4000 K and $P$ = 0-10 GPa show a continuous increase in cell volume (Supplementary Figure 9), indicating that the bulk crystal tends to sublimate into small molecules and single atoms. This evaporation is expected at low pressures and high temperatures. In our two-phase coexistence simulation at a low pressure of 10 GPa, the decomposed phase presents significant deformation and stress, further indicating instability of the bulk structure under these conditions.
@@
Figure 6: Phase diagram of SiC. Results from our MD simulations are denoted in big squares, and the inferred phase boundaries are drawn with black solid lines (interpolated) and dash lines (extrapolated). Blue: B3 (ZB), green: B1 (RS), orange: decomposed Si + C, red: homogeneous liquid, brown: gas. The background colors indicate the stability regions of different phases inferred by our simulations. Experimental reports of congruent melting (blue-red): diamond [12], triangle [13]. Incongruent melting (blue-orange): diamond [11], triangle [35], circle [10], plus [41], cross [8], upside-down triangle [9]. B3-B1 transition (blue-green): cross [42], triangle [43], diamond [44], circle [45], plus [46], upside-down triangle [47].
The transition temperatures that we observe between B3 and decomposed Si + C at high pressures are significantly higher than those reported in a previous experiment by Daviau et al. (blue-orange circles) [10]. This discrepancy may be attributed to the challenges in high-pressure measurements and the likely presence of significant temperature gradients and transients in the diamond anvil setup (see more details in the Discussions section). Nevertheless, at pressures around 10 GPa, our transition temperature is consistent with experimental results by Togaya et al. (blue-orange upside-down triangles) [9].
@@
The high-temperature phase boundary is close to the melting point of graphite and diamond carbon, indicating that the C clusters dissolve at around 4000 K and SiC becomes a homogeneous liquid mixture. As depicted in Supplementary Figure 1, the Si-C peak indicates that the Si and C are mixed well, and the complete phase separation does not occur, as opposed to the RDF shown in Fig. 2.
4 Discussions
@@
Our MLFF, based on Gaussian process regression [26], provides uncertainty estimations for its predictions. To ensure the fidelity of our large-scale MD simulations, we assess whether high-uncertainty configurations are involved. A sufficient training dataset covering a variety of configurations results in low uncertainties during simulations, as the model is familiar with the encountered atomic structures. Conversely, high model uncertainty implies the MD has proceeded to configurations not covered by the training data, potentially reducing prediction reliability.
Therefore, we evaluate the model uncertainty on snapshots from the decomposition MD simulations. As shown in Supplementary Figure 11, the model uncertainty remains well below the acquisition criteria of our force field active learning workflow. This indicates that our collected training data set sufficiently covers the configuration space, ensuring high confidence level in the model predictions.
@@
While numerous experiments have confirmed the occurrence of incongruent melting or decomposition in SiC, some studies have reported congruent melting. Considering the slow kinetics of SiC decomposition and phase transformation, the grain size can influence the process. For instance, in experiments reporting congruent melting, large single crystalline grains (between 150 $\mu$m and 3 mm) are used, which likely hinders the transition [13].
@@
While our MD simulations confirm incongruent melting at various pressures, qualitatively consistent with the experiment [10], our obtained transition temperatures at high pressures are significantly higher than the reported measurements around 2000 K. This discrepancy may be attributed to several factors in experimental setups: (1) Since the temperature is measured via the thermal emission from the surface, the approach could underestimate the core temperature. (2) The heat distribution in the sample could be inhomogeneous, such that overheating in certain regions might facilitate the decomposition at lower apparent temperatures. Although there are no other experimental reports at such high pressures, our simulation is in good agreement with the transition temperature at a lower pressure of 10 GPa found in another experiment [9]. We also note that in other applications, such as physical deposition vapor growth of SiC, a temperature of $\sim$ 2500 K is used at $\sim$ 0 GPa, demonstrating SiC’s stability at temperatures higher than 2000 K [48]. This observation further supports the plausibility of our higher predicted transition temperatures at elevated pressures.
@@
The two-phase coexistence MD simulation creates an interface between two different phases and thereby overcomes the nucleation barrier. It is possible that the defect concentration influences the phase transition temperature. Even state of the art epitaxially grown SiC wafers have high defect densities [48], implying that non-optimized powered SiC contains large amounts of defects. Therefore, we performed additional simulations, in which we introduce different point defect concentrations in the crystal. As shown in Supplementary Figure 7, the concentration of defects does not have a significant influence on the transition temperature converged by the two-phase MD simulation. Macroscopic defects such as dislocations and phase boundaries could have a greater impact than point defects but are out of scope of this study.
@@
While some previous computational studies using empirical and neural network potentials obtain amorphous SiC through melting and quenching [20, 21], our simulations show that the melting and quenching process results in the decomposition of the SiC. Instead, techniques such as irradiation are required for the amorphization of SiC, which is consistent with most of the experiments [22, 23].
@@
Building upon our findings on SiC decomposition, we now explore possible mechanisms underlying this process. The two primary mechanisms to consider are nucleation and spinodal decomposition. To investigate spinodal decomposition, we calculated the C-C spatial correlation function from local concentrations (Supplementary Figure 10). Notably, towards the end of cooling process, particularly below 3400 K, a peak appears at approximately 40 Å. This implies periodic fluctuations in the carbon concentration, indicative of a wavelength of the phase separation mode [49, 50]. This observation aligns with the spinodal decomposition mechanism reported in alloy systems [49, 50] and is consistent with a recent tight-binding MD study of SiC under tensile stresses [51]. To analyze the nucleation of solid carbon clusters during decomposition, we tracked the evolution of diamond and graphite clusters. Fig.7 illustrates the various stages of carbon cluster development, highlighting their transformation into either diamond or graphite phases. We applied cluster analysis in OVITO [37] with a 1.55 Å cutoff - approximately the bond length in graphite and diamond - to identify clusters. The formation of different phases were analyzed by the "identify diamond structure" method [52] for diamond, and polyhedral template matching [36] for graphite. We track the evolution of clusters by identifying those with the largest overlap with the final structures at each time step. As shown in Fig.7, in earlier stages, small nuclei lack clear lattice symmetry or phase identity. For graphite, the hexagonal structure can be traced back to smaller clusters. However, tracking becomes challenging for clusters below 100 atoms, as dynamic rearrangements reduce the overlap ratio quickly over time for both diamond and graphite structures.
@@
In conclusion, our machine learning molecular dynamics study confirms the occurrence of incongruent melting in SiC, with the decomposition of SiC into Si and C being observed at a temperature range of 3200 - 3600 K at high pressures. We also identified a transition at higher temperatures where the decomposed phase transforms into a homogeneous liquid phase. The highly efficient and accurate MLFF enabled previously intractable modeling and simulation of the decomposition process, overcoming limitations of both ab initio or empirical force fields in terms of computational cost and accuracy. Our MD simulations provide atomic-level insights, revealing nucleation mechanisms and process details inaccessible in experiments. These findings support the existence of incongruent melting at high pressure, clarifying controversial experimental observations. In summary, our study provides a comprehensive understanding of SiC melting and decomposition behavior, along with a complete phase diagram, reconciling and extending previous experimental and theoretical work in this field.

### v2

@@ 2.4 Low-temperature phase boundary: crystal $\leftrightarrow$ decomposed Si+C
Figure 5: Analysis of two-phase coexistence simulations. Panels (left to right) show: fraction of zinc-blende (rock-salt) structures, largest C cluster size (number of atoms), temperature convergence at the NPHNPH stage, over simulation time. Top: 30 GPa. Bottom: 90 GPa.
@@
At the fixed pressure of 30 GPa, during the heating stage (0.5–1 ns), the heated half of the supercell becomes amorphous, and the overall fraction of zinc-blende local structures in the supercell decreases to 0.5. The C clusters begin to form during the cooling stage (1–1.5 ns) and continue growing in size at 1.5–2 ns as demonstrated in the second column of Fig. 5. In the NPH simulation stage (1.5–2 ns), both B3 crystal and decomposed Si + C phases coexist, with relatively different equilibrated fractions dependent on the initial temperature. Meanwhile, multiple simulations with different initial temperatures converge to approximately the same temperature during the NPH stage, which is identified as the transition temperature (shown in the third column of Fig. 5).
At 90 GPa, where the stable phase at low temperature is rock salt (B1), a similar procedure starting from B1 phase yields a transition temperature around 4500 K. Notably, the carbon cluster sizes formed at 90 GPa are much smaller than those at 30 GPa, leading us to conjecture that at sufficiently high pressures, the decomposed Si+C phase between the crystalline and liquid phases may not be present. Additional coexistence simulations were performed for the B3 phase (at 10, 20, 45, 60, 70, and 80 GPa) and the B1 phase (at 85 and 90 GPa). In addition, we investigated finite-size effects by performing significantly larger simulations (4x and 8x the original size). These results and the details for determining the final phase transition temperatures are presented in the Supplementary Section 7.3, 7.4, and 7.6.
@@ 2.5 Phase diagram
Finally, we combine the results obtained from the MD simulations of the SiC phase transitions into a complete pressure-temperature phase diagram. Fig. 6 presents the $P$-$T$ diagram, illustrating the B3 $\leftrightarrow$ B1, B3 $\leftrightarrow$ gas, B3 $\leftrightarrow$ Si + C, B1 $\leftrightarrow$ Si + C, and Si + C $\leftrightarrow$ SiC (liquid) transitions. For completeness, we briefly describe the simulations of B3 $\leftrightarrow$ B1 transition and B3 $\leftrightarrow$ gas below, although these are not the primary focus of this work.
@@
The B3 $\leftrightarrow$ B1 transition pressures at various temperatures are determined using the same methods as in our previous work [32]. Starting with the B3 crystal, we increase pressure at a fixed temperature to generate a configuration containing both phases. Then, we perform constant-pressure MD simulations on these coexistence structures at different pressures to identify the transition point, characterized by the stable coexistence of two phases remains. The pressure scan results are shown in the Supplementary Section 8.1. The B3 $\leftrightarrow$ B1 transition pressures identified from our simulations (along pink-blue boundary in Fig. 6) are close to experimental measurements with a deviation of less than 20 GPa.
At zero or low pressure, the SiC B3 crystal sublimates into gas phase instead of melting, which only occurs under high pressure. Our $NPT$ MD simulations at $P$ = 0–15 GPa and $T$ = 3900–4500 K reveal a qualitative pressure threshold between 5 and 10 GPa: at lower pressures (0, 1, and 5 GPa), sustained volumetric expansion above 4000–4200 K indicates progressive sublimation to a gas/vapor phase (details in Supplementary Section 8.2). While our simulations identify this sublimation threshold, most experimental literature at low pressure focuses on sublimation growth conditions rather than on establishing the equilibrium solid-gas phase boundary. An earlier study [33] reported a sublimation temperature-pressure relationship, but it did not distinguish between the specific solid phases of SiC.
@@
Figure 6: Phase diagram of SiC. Results from our MD simulations are denoted in big squares, and the inferred phase boundaries are drawn with black solid lines (interpolated) and dash lines (extrapolated). Blue: B3 (ZB), pink: B1 (RS), orange: decomposed Si + C, red: homogeneous liquid, brown: gas. The background colors indicate the stability regions of different phases inferred by our simulations. Experimental reports of congruent melting (blue-red symbols): hexagons [13], rotated hexagons [14]. Incongruent melting (blue-orange symbols): diamonds [10], triangles [11], pluses [12], crosses [7], left-pointing triangles [8], circles [9]. B3-B1 transition (blue-pink): triangles [34], crosses [35], diamonds [36], circles [37], pluses [38], inverted triangles [39].
Regarding the transition between condensed phases, there is a qualitative disagreement among experiments: Hall [13] and Sokolov [14] report congruent melting, whereas Daviau et al. [9] and five earlier independent studies (Bhaumik et al. [10, 11], Dolloff [12], Togaya [8], Ekimov et al. [7]) observe incongruent decomposition—consistent with our simulations. Quantitatively, at 10 GPa, our predicted decomposition temperature falls within the cluster formed by those five independent measurements, supporting our placement of the boundary there. At higher pressures, however, our values are substantially above the laser-heated DAC series from Daviau et al. [9] (blue-orange circles). This laser-heated DAC dataset shows several inconsistencies. When extrapolated downward, it sits discontinuously below the lower-pressure data. It also shows internal inconsistency with a separate study by the same group [34] (blue triangles). Furthermore, it conflicts with other B3-B1 studies [35, 36] (blue cross and pink-blue diamonds) that report stable B3-SiC in a region where Daviau et al. claimed decomposition occurs. Together with the broader scatter among datasets, this indicates that the high-pressure decomposition boundary remains experimentally underconstrained. Potential sources for these discrepancies, such as temperature gradients and transients in the DAC environment, highlight the need for further high-pressure measurements to resolve these conflicts.
@@
The high-temperature phase boundary is close to the melting point of graphite and diamond carbon, indicating that the C clusters dissolve at around 4000 K and SiC becomes a homogeneous liquid mixture. As depicted in Supplementary Fig. S3, the Si-C peak indicates that the Si and C are mixed well, and the complete phase separation does not occur, as opposed to the RDF shown in Fig. 2.
3 Discussion
Uncertainty aware simulations
@@ 3 Discussion
Our MLFF, based on Gaussian process regression [40], provides uncertainty estimations for its predictions. To ensure the fidelity of our large-scale MD simulations, we assess whether high-uncertainty configurations are involved. A sufficient training dataset covering a variety of configurations results in low uncertainties during simulations, as the model is familiar with the encountered atomic structures. Conversely, high model uncertainty implies the MD has proceeded to configurations not covered by the training data, potentially reducing prediction reliability. Therefore, we evaluate the model uncertainty on snapshots from the decomposition MD simulations. As shown in Supplementary Figure S16, the model uncertainty remains well below the acquisition criteria of our force field active learning workflow. This indicates that our collected training data set sufficiently covers the configuration space, ensuring high confidence level in the model predictions.
Heating-cooling protocol
@@
Grain size effects on congruent melting observations
[…]
Limitations of simulations and experiments
While our MD simulations confirm incongruent melting at various pressures, qualitatively consistent with Daviau et al. [9], our obtained transition temperatures at high pressures are significantly higher than the reported measurements around 2000 K. One limitation of our MLFF is that it is trained primarily on high-temperature/high-pressure data, focusing on bulk solid-solid and solid-liquid transitions, but not optimized for simulating surfaces or capturing detailed effects across a wide range of complex crystalline defect ensembles. The expressive power of MLFF and the accuracy of DFT can contribute to the discrepancy of our simulation results with experiments. More expressive and expensive MLFF models such as the equivariant neural works [41, 42] can be used to potentially improve the accuracy.
@@
On the other hand, the discrepancy may also be attributed to several factors in experimental setups: (1) Since the temperature is measured via the thermal emission from the surface, the approach could underestimate the core temperature. (2) The heat distribution in the sample could be inhomogeneous, such that overheating in certain regions might facilitate the decomposition at lower apparent temperatures. Although there are no other experimental reports at such high pressures, our simulation is in good agreement with the transition temperature at a lower pressure of 10 GPa found in five other experiments (Bhaumik et al. [10, 11], Dolloff [12], Togaya [8], and Ekimov et al. [7]). We also note that in other applications, such as physical deposition vapor growth of SiC, a temperature of $\sim$ 2500 K is used at $\sim$ 0 GPa, demonstrating SiC’s stability at temperatures higher than 2000 K [43]. This observation further supports the plausibility of our higher predicted transition temperatures at elevated pressures.
Effect of point defects and surfaces
@@
The two-phase coexistence MD simulation creates an interface between two different phases and thereby overcomes the nucleation barrier. It is possible that the defect concentration influences the phase transition temperature. Even state-of-the-art epitaxially grown SiC wafers have high defect densities [43]. Therefore, we performed additional simulations, in which we introduce different point defect concentrations in the crystal. As shown in Supplementary Section 7.5, the concentration of defects does not have a significant influence on the transition temperature converged by the two-phase MD simulation. Macroscopic defects such as dislocations and phase boundaries could have a greater impact than point defects but are out of scope of this study. Besides point defects, we further note that free surfaces or high surface-to-volume ratios, absent from our periodic bulk simulations, can modify local coordination, provide heterogeneous nucleation sites, and shift the decomposition/melting temperature, and thus may contribute to remaining experiment–simulation discrepancies.
Amorphous SiC formation
@@
While some previous computational studies using empirical and neural network potentials obtain amorphous SiC through melting and quenching [21, 22], our simulations show that the melting and quenching process results in the decomposition of the SiC. Instead, techniques such as irradiation are required for the amorphization of SiC, which is consistent with most of the experiments [44].
Decomposition mechanisms: nucleation and spinodal decomposition
@@
Building upon our findings on SiC decomposition, we now explore possible mechanisms underlying this process. The two primary mechanisms to consider are nucleation and spinodal decomposition. To investigate spinodal decomposition, we calculated the C-C spatial correlation function from local concentrations (Supplementary Section 6.5). Notably, towards the end of cooling process, particularly below 3400 K, a peak appears at approximately 40 Å. This implies periodic fluctuations in the carbon concentration, indicative of a wavelength of the phase separation mode [45, 46]. This observation aligns with the spinodal decomposition mechanism reported in alloy systems [45, 46] and is consistent with a recent tight-binding MD study of SiC under tensile stresses [47]. To analyze the nucleation of solid carbon clusters during decomposition, we tracked the evolution of diamond and graphite clusters. Fig. 7 illustrates the various stages of carbon cluster development, highlighting their transformation into either diamond or graphite phases. We applied cluster analysis in OVITO [29] with a 1.55 Å cutoff — approximately the bond length in graphite and diamond — to identify clusters. The formation of different phases were analyzed by the “identify diamond structure” method [48] for diamond, and polyhedral template matching [28] for graphite. We track the evolution of clusters by identifying those with the largest overlap with the final structures at each time step. As shown in Fig. 7, in earlier stages, small nuclei lack clear lattice symmetry or phase identity. For graphite, the hexagonal structure can be traced back to smaller clusters. However, tracking becomes challenging for clusters below 100 atoms, as dynamic rearrangements reduce the overlap ratio quickly over time for both diamond and graphite structures.
@@ Conclusions
In conclusion, our machine learning molecular dynamics study confirms the occurrence of incongruent melting in SiC, with the decomposition of SiC into Si and C being observed at a temperature range of 3000 – 3600 K at high pressures. We also identified a transition at higher temperatures where the decomposed phase transforms into a homogeneous liquid phase. The highly efficient and accurate MLFF enabled previously intractable modeling and simulation of the decomposition process, overcoming limitations of both ab initio or empirical force fields in terms of computational cost and accuracy. Our MD simulations provide atomic-level insights, revealing nucleation mechanisms and process details inaccessible in experiments. These findings support the existence of incongruent melting at high pressure, clarifying controversial experimental observations. In summary, our study provides a comprehensive understanding of SiC melting and decomposition behavior, along with a complete phase diagram, reconciling and extending previous experimental and theoretical work in this field.
4 Methods
Machine learning force field
Accurately modeling the melting and decomposition of SiC at high temperature and pressure requires an interatomic potential that captures complex bonding environments, including phase separation. Traditional empirical force fields can reproduce certain crystalline and shock-response properties. For example, Tersoff [19] and Vashishta [17] potentials have been applied in planar and impact shock studies of SiC [49, 50, 51]. However, the concurrent bond breaking, chemical partitioning, and long-timescale phase coexistence needed to delineate the high-$P$ / high-$T$ decomposition boundary remains challenging for such potentials. While AIMD [15, 16] provides accurate atomic forces, its high computational cost limits simulations to small system sizes and short timescales. To overcome these challenges, we employ a MLFF trained on DFT calculations, such that the atomic interactions can be modeled with high accuracy, while remaining cost effective for large-scale simulations.
Among general MLFF frameworks [52, 53, 41, 42], FLARE was selected because: (i) Bayesian uncertainties focus acquisition on novel crystal, liquid, and early segregation environments, enabling highly efficient collection of non-redundant data; (ii) sparse GP to polynomial mapping yields high simulation efficiency at the $10^{4}$$10^{6}$ atom scale needed for multi-temperature/pressure coexistence runs; (iii) reuse of a previously validated SiC setup [32] ensures methodological continuity. More expressive deep equivariant models could lower force errors at substantially higher cost, a trade-off we reserve for future refinements.
Our MLFF is built using the FLARE framework [26, 32], which is based on a sparse Gaussian process (SGP) model [40] with atomic cluster expansion descriptors [54]. This model predicts forces while providing uncertainty quantification.
Bayesian active learning
To efficiently generate training data, we employ Bayesian active learning, which systematically selects new atomic configurations where the MLFF is most uncertain. The workflow follows these steps: (1) Initial Training: The MLFF is first trained on a small set of DFT-calculated structures, including perfect and slightly perturbed crystal structures. (2) Uncertainty-Guided MD: The MLFF runs MD simulations, predicting atomic forces while simultaneously estimating its uncertainty. (3) Selective DFT Calculations: If the model’s uncertainty exceeds a threshold, new configurations are labeled using DFT. Otherwise, continue with MD. (4) Iterative Refinement: The MLFF is retrained with the newly acquired DFT data, progressively improving its accuracy. This adaptive learning process ensures that the force field is only trained where necessary, avoiding redundant DFT calculations while capturing the relevant phase space. Once the active learning is complete, we map the SGP into an equivalent but more computationally efficient polynomial model [40, 32], which serves as the final MLFF for large-scale production simulations in LAMMPS [55]. The DFT calculations are performed using VASP [56] with the PBE exchange correlation functional [57], following the computational settings in our previous works [58, 32].
The initial training set contains SiC polymorphs data from our previous study [32], encompassing the hexagonal 2H, 4H, and 6H polytypes, B3 (zinc blende or 3C) and the high-pressure B1 (rock salt) phases. We then employ the FLARE Bayesian active learning workflow to systematically sample data that captures melting and decomposition behavior SiC across a wide pressure range. Active learning MD simulations are performed for a 64-atom SiC supercell at 0, 30, 60, 90, and 120 GPa over a temperature range of 2000–6000 K until melting is observed. In addition, active learning for a 512-atom SiC supercell is conducted at 5000 K across the same pressure range. To enhance our data coverage for the decomposition products, additional data is collected for pure carbon (diamond and graphite phases, 2000–4000 K) and pure silicon (1000–3000 K) across all pressure conditions.
The final trained MLFF is then used to investigate SiC melting and decomposition under high temperature and pressure conditions through two-phase coexistence simulations, which allowed for the accurate determination of phase transition temperatures.
Large-scale molecular dynamics simulation
The large-scale MD simulations are conducted using GPU-accelerated LAMMPS [55] with Kokkos parallelization [59] and the FLARE MLFF pairstyle, which has demonstrated exceptional performance in billion-atom catalytic simulations [60]. During the melting simulations, the NPT ensemble with anisotropic (aniso) control of cell dimensions is employed, allowing independent scaling along the $x$, $y$ and $z$ directions through stress components $P_{xx}$, $P_{yy}$, and $P_{zz}$. Phase separation exhibits a tendency to minimize the interface between decomposed phases, necessitating anisotropic elongation of the supercell, as illustrated in Fig. 2.
5 Code Availability
The scripts are available on Github: https://github.com/YuuuXie/SiC_MLMD_phase_diagram.